# Generalized symmetry in non-Hermitian systems

### v1

@@ V Generalized unitary symmetry
In this section, we further explore the deep contrast between the consistent formalism of non-Hermitian quantum mechanics, embodied by BQM and metric formalisms, and the oft-used no-jump approximation valid for dissipative systems. Specifically, we focus on aspects of symmetry. We show that the consistent formulation of non-Hermitian quantum mechanics leads to a generalization of the conventional notions of symmetry associated with the Hermitian case. Using a toy model, we show that the generalization of a discrete symmetry in the model leads to important symmetry constraints for observables. We contrast this with the no-jump formalism where such constraints are absent.
A first point of significance is that the left and right eigenvectors of a non-Hermitian Hamiltonian do not coincide, unlike the Hermitian case. Therefore, the construction of symmetry operators should be generalized to incorporate this aspect. The biorthogonal basis provides a natural path to this generalization. The spectral decomposition for a general non-Hermitian Hamiltonian, parameterized by a parameter $k$, acting on a finite-dimensional Hilbert space reads as follows:
@@
We note that a similar notion also arises naturally within the metric framework. To this end, we can define a unitary operator similar to Eq. (27) using the spectrum of the Hermitian mapping $h_{k}$, given by Eq. (8). Writing $h_{k}=\sum_{n}e_{k,n}\lvert k,n\rangle\langle k,n\lvert$, the symmetry is defined as
$U_{k}=\sum_{n}\lvert-k,n\rangle\langle k,n\lvert.$ (29)
It then follows that the even parity of the spectrum $e_{k}=e_{-k}$ translates directly to the symmetry $U_{k}h_{k}U_{k}^{\dagger}=h_{-k}$. This shows that the generalized unitary symmetry is already embedded in the metric framework, and we can access it directly in the observables without the need to consider BQM.
@@
where $k$ denotes the momentum, as was considered in Ref. [46]. This Hamiltonian does not necessarily have to satisfy Eq. (14). In general, the coefficients $\gamma$ and $\Delta$ in Eq. (30) can be time dependent, but we omit the time dependence hereafter for the ease of notation.
Ref. [46] showed that the metric dynamics led to qualitatively different symmetry structures of spin expectation values. This was shown to have important physical consequences compared to the straightforward renormalization of the no-jump formalism. Here we show that this symmetry structure is a direct consequence of the generalized symmetry discussed above. To this end, we consider the left and right instantaneous eigenstates of Eq (30), which are denoted $\lvert k,\pm\rangle_{L/R}$, as in Eq. (1). They are normalized such that they obey Eqs. (2) and (3). The corresponding eigenvalues are $E_{k}=\pm\sqrt{k^{2}-\gamma^{2}+\Delta^{2}}$, such that they satisfy the requirement for Eq. (27). The spectrum is real in the parameter regime $k^{2}+\Delta^{2}\geq\gamma^{2}$ where $\mathcal{PT}$-symmetry is not spontaneously broken, and form complex conjugate pairs otherwise. The exact forms of $U_{L,k}$ and $U_{R,k}$ for our model are given by $U_{R,k}=U_{R,k}^{\dagger}=\frac{1}{\gamma+k}(\gamma\mathbbm{1}-k\sigma_{z})$ and $U_{L,k}=U_{L,k}^{\dagger}=\frac{1}{\gamma-k}(\gamma\mathbbm{1}+k\sigma_{z})$, such that Eq. (28) is satisfied. For this model, Eq. (29) greatly simplifies, giving $U_{k}=\sigma_{z}$. This implies the symmetry $\sigma_{z}h_{k}(t)\sigma_{z}=h_{-k}(t)$.
@@
Introducing the pure-state density matrix of the time-evolved state $\hat{\Psi}_{k}(t)\equiv\lvert\Psi(t)\rangle_{k~{}k}{\langle\Psi(t)\lvert}$, it is easy to verify the symmetry properties of the spin expectation values with respect to $k\rightarrow-k$. The density matrix is given by
@@ VI Conclusion
Our findings indicate that when non-Hermiticity arises from dissipation in the limit where quantum jumps can be neglected, the normalization method is more appropriate, especially in the $\mathcal{PT}$-broken regime. This method closely reproduces the qualitative behavior of the full master equation in this regime. Conversely, when non-Hermiticity is directly engineered in the system, such as by Naimark dilation, it is essential to consider the metric dynamics for a consistent quantum treatment. This approach can reveal richer physics, including the manifestation of a generalized symmetry. Notably, our findings align with those of Ref. [54], which highlights starkly different behaviors obtained by BQM and explicit normalization in the context of dynamical phase transitions in non-Hermitian systems.
Our work paves the way for deeper explorations of truly non-Hermitian Hamiltonians, where the physics of the metric and the concept of generalized symmetries could potentially be harnessed for exotic behaviors. An exciting future direction is the extension of this framework to include other discrete and continuous symmetries. This paves the way for the exploration of new paradigms for quantum criticality, spontaneous symmetry breaking and plausibly new classification schemes based on generalized symmetries for topological states.
@@ Acknowledgements.
This research was funded by the Swiss National Science Foundation (SNSF) grant number 200021 207537, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC2181/1-390900948 (the Heidelberg STRUCTURES Excellence Cluster), the Swiss State Secretariat for Education, Research and Innovation (SERI), and by the Swedish Research Council (2018-00313) and Knut and Alice Wallenberg Foundation (KAW) via the project Dynamic Quantum Matter (2019.0068). The authors would like to thank G. M. Graf for numerous fruitful discussions.

### v2

@@ V Generalized unitary symmetry
In this section, we further explore the deep contrast between the consistent formalism of non-Hermitian quantum mechanics, embodied by BQM and metric formalisms, and the oft-used no-jump approximation valid for dissipative systems. Specifically, we focus on aspects of symmetry. We show that the consistent formulation of non-Hermitian quantum mechanics leads to a generalization of the conventional notions of symmetry associated with the Hermitian case. This generalization is made by taking into account the left and right eigenvectors in its construction. Using a toy model, we show that the generalization of a discrete symmetry in the model leads to important symmetry constraints for observables.
The construction of generalized symmetry operators also sheds light on the results of Ref. [50]. Using the metric formalism, it was shown that the spin expectation values possess fixed parity with respect to momentum. In contrast, the symmetry of the spin expectation values only emerged in the asymptotic limit when physical quantities are simply divided by the time-dependent norm. This division by the time-dependent norm is equivalent to the no-jump approximation in dissipative systems, as shown by Eq. (13). This presence (or absence) of the symmetry of the observables at finite times is illustrated in Fig. 2. The symmetry constraints in the metric framework can be explained by using the generalized symmetry operators. Therefore, we focus on the discrete symmetry of $k\rightarrow-k$ in this section to demonstrate the concept of generalized symmetry operators, where $k$ denotes momentum.
Figure 2: The spin expectation value $\langle\sigma_{z}(t)\rangle_{k}$ at different times, calculated using (a) the metric framework, Eq. (10) and (b) the no-jump approximation (division by norm), Eq. (13). The system is evolved using the non-Hermitian Hamiltonian given in Eq. (32), with $\gamma=1$ and $\Delta(t)=t$. The initial state at $t\rightarrow-\infty$ is $(1,0)^{T}$. Fig. (a) shows the even $k\rightarrow-k$ symmetry of $\langle\sigma_{z}\rangle_{k}$ within the metric framework. This even symmetry is present at all times. Fig. (b) shows that within the no-jump approximation, the odd $k\rightarrow-k$ symmetry of $\langle\sigma_{z}\rangle_{k}$ only arises in the asymptotic limit when $t\rightarrow\infty$. For finite times, there is no definite parity in the momentum $k$.
A first point of significance in a non-Hermitian system is that the left and right eigenvectors of the Hamiltonian do not coincide, unlike the Hermitian case. Therefore, the construction of symmetry operators should be generalized to incorporate this aspect. The biorthogonal basis provides a natural path to this generalization. The spectral decomposition for a general non-Hermitian Hamiltonian, parameterized by a parameter $k$, acting on a finite-dimensional Hilbert space reads as follows:
@@
We note that this left-right construction of symmetry operators is not restricted to unitary symmetries. In particular, this construction is also applicable to operations involving complex conjugation or transpose, which are shown to be different for non-Hermitian Hamiltonians [60]. For example, we can demonstrate this by considering the case of the particle-hole symmetry (PHS) operators, $PHS$ and $PHS^{\dagger}$ considered in Ref. [60]. We take the $PHS:=P$ operator defined in Ref. [60] to be the right symmetry operator $P_{R}$, such that
$H=-P_{R}H^{T}P_{R}^{-1},\quad\quad\quad H=-P_{R}^{\dagger}H^{\ast}(P_{R}^{\dagger})^{-1}.$ (29)
The corresponding left symmetry operator $P_{L}$ can then be defined as
$H^{\dagger}=-P_{L}(H^{\dagger})^{T}P_{L}^{-1},\quad\quad\quad H^{\dagger}=-P_{L}^{\dagger}(H^{\dagger})^{\ast}(P_{L}^{\dagger})^{-1}.$ (30)
We now compare Eq. (30) with the result obtained from the Hermitian conjugation of Eq. (29). This gives $P_{L}^{-1}=P_{R}^{\dagger}$ and $P_{R}^{-1}=P_{L}^{\dagger}$. Therefore, given the right PHS symmetry operators, we can find the corresponding left symmetry operators as described above.
A similar notion also arises naturally within the metric framework. To this end, we can define a unitary operator similar to Eq. (27) using the spectrum of the Hermitian mapping $h_{k}$, given by Eq. (8). Writing $h_{k}=\sum_{n}e_{k,n}\lvert k,n\rangle\langle k,n\lvert$, the symmetry is defined as
$V_{k}=\sum_{n}\lvert-k,n\rangle\langle k,n\lvert.$ (31)
It then follows that the even parity of the spectrum $e_{k}=e_{-k}$ translates directly to the symmetry $V_{k}h_{k}V_{k}^{\dagger}=h_{-k}$. As stated in Sec. II.2, $h_{k}$ and $H_{k}$ are equivalent representations of the same non-Hermitian system. Accordingly, the observables calculated using the evolution with $h_{k}$ and $H_{k}$ are the same, as shown in Eq. (10). Therefore, the observable calculated within the metric framework incorporates the symmetry imposed by $V_{k}$, and thus $U_{R,k}$ and $U_{L,k}$. This shows that the notion of generalized symmetry is already embedded in the metric framework, and we can access it directly in the observables without the need to consider BQM.
@@
where $k$ denotes the momentum, as was considered in Ref. [50]. This Hamiltonian does not necessarily have to satisfy Eq. (14). In general, the coefficients $\gamma$ and $\Delta$ in Eq. (32) can be time dependent, in which case the symmetry operators defined in Eq. (27) should be constructed using the instantaneous eigenstates of the Hamiltonian. For a time-dependent Hamiltonian, the symmetry operators are thus time-dependent in general.
As stated earlier, Ref. [50] showed that the metric dynamics led to qualitatively different symmetry structures of spin expectation values. This was shown to have important physical consequences compared to the straightforward division by norm of the no-jump formalism. Here we show that this symmetry structure is a direct consequence of the generalized symmetry discussed above. To this end, we construct the symmetry operators for the Hamiltonian considered in Ref. [50], which corresponds to setting $\Delta(t)=Ft$ and $\gamma(t)=\gamma$ in Eq. (32), where $F$ and $\gamma$ are constant. The corresponding eigenvalues of the Hamiltonian are $E_{k}=\pm\sqrt{k^{2}-\gamma^{2}+\Delta^{2}}$, such that they satisfy the requirement for Eq. (27). The spectrum is real in the parameter regime $k^{2}+\Delta^{2}\geq\gamma^{2}$ where $\mathcal{PT}$-symmetry is not spontaneously broken, and form complex conjugate pairs otherwise.
Using the instantaneous eigenstates of the Hamiltonian, we construct the symmetry operators $U_{L,k}$ and $U_{R,k}$ as in Eq. (27). Their exact forms are given by $U_{R,k}=U_{R,k}^{\dagger}=\frac{1}{\gamma+k}(\gamma\mathbbm{1}-k\sigma_{z})$ and $U_{L,k}=U_{L,k}^{\dagger}=\frac{1}{\gamma-k}(\gamma\mathbbm{1}+k\sigma_{z})$, such that Eq. (28) is satisfied. For this model, Eq. (31) greatly simplifies, giving $V_{k}=\sigma_{z}$. This implies the symmetry $\sigma_{z}h_{k}(t)\sigma_{z}=h_{-k}(t)$.
@@
Introducing the pure-state density matrix of the time-evolved state $\hat{\Psi}_{k}(t)\equiv\lvert\Psi(t)\rangle_{k~{}k}{\langle\Psi(t)\lvert}$, it is easy to verify the symmetry properties of all 3 spin expectation values with respect to $k\rightarrow-k$. The density matrix is given by
@@ VI Conclusion
Our findings indicate that when non-Hermiticity arises from dissipation in the limit where quantum jumps can be neglected, it is more appropriate to simply normalize physical quantities by dividing by the norm of the time-evolved state, especially in the $\mathcal{PT}$-broken regime. This method closely reproduces the qualitative behavior of the full master equation in this regime. Conversely, when non-Hermiticity is directly engineered in the system, such as by Naimark dilation, it is essential to consider the metric dynamics for a consistent quantum treatment. This approach can reveal richer physics, including the manifestation of a generalized symmetry. Notably, our findings align with those of Refs. [61, 62], which highlights starkly different behaviors obtained by BQM and division by the norm in the context of dynamical phase transitions in non-Hermitian systems.
Our work paves the way for deeper explorations of truly non-Hermitian Hamiltonians, where the physics of the metric and the concept of generalized symmetries could potentially be harnessed for exotic behaviors. An exciting future direction is the extension of this framework to include other discrete and continuous symmetries. Although we would need other parameters to index different types of symmetries, it would still be necessary to include the left and right eigenstates in the construction of the symmetry operators. This paves the way for the exploration of new paradigms for quantum criticality, spontaneous symmetry breaking and plausibly new classification schemes based on generalized symmetries for topological states.
@@ Acknowledgements.
This research was funded by the Swiss National Science Foundation (SNSF) grant numbers 200021 207537 and 200020 184739, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC2181/1-390900948 (the Heidelberg STRUCTURES Excellence Cluster), the Swiss State Secretariat for Education, Research and Innovation (SERI), and by the Swedish Research Council (2018-00313) and Knut and Alice Wallenberg Foundation (KAW) via the project Dynamic Quantum Matter (2019.0068). The authors would like to thank G. M. Graf for numerous fruitful discussions.